OpenStudy (anonymous):

Medal~What is the slope of the line? A. 0 B. 5 C. 11 D. Undefined

OpenStudy (anonymous):

@elementwielder @rock_mit182 @Stormswan @davidmack @thomaster

OpenStudy (anonymous):

@superdavesuper

OpenStudy (anonymous):

i think its 0

OpenStudy (superdavesuper):

slope of a line = (y1-y2)/(x1-x2) where (x1, y1) and (x2, y2) are two points on the line ok?

OpenStudy (anonymous):

yea...

OpenStudy (superdavesuper):

so put the two points in the eqn....what will u get?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

2-2=0

OpenStudy (superdavesuper):

yes 2-2=0 so the slope = 5/0 which is.....?

OpenStudy (anonymous):

5 divided by 0 = cant divide by 0

OpenStudy (superdavesuper):

also correct so the right answer should be.....

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

@superdavesuper

OpenStudy (superdavesuper):

nope u said it "cant divide by 0" so it is undefined.

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

a slope where y=0 is undefined but if x is a straight line then the slope is =y-intercept